1. How many permutation of 5 different digits can be arrange from the numbers 0 to 20
2. How many permutation of 3 different letters can be arrange from the word “NOTEBOOK” (repetition is not allowed)
3. A password is consists of 4 characters, the two characters is made of alpha characters from A-E and the other two characters is made from F-G. How many arrangement of character can be made?
4. How many words can be made from the word “KEYBOARD” with any length?

Answer by Edwin McCravy(20059)


1. How many permutation of 5 different digits can be arrange from the numbers 0 to 20
There are 21 elements in this set

{0,1,2,3,4,5,6,7,8,9,10,1,12,13,14,15,16,17,18,19,20}

P(21,5) = 21·20·19·18·17 = 2441880

2. How many permutation of 3 different letters can be arrange from the word “NOTEBOOK” (repetition is not allowed)
That's the set {N,O,T,E,B,K} which contains 6 elements.

P(6,3) = 6·5·4 = 120.

3. A password consists of 4 characters, the two characters is made of alpha characters from A-E and the other two characters is made from F-G. How many arrangement of character can be made?
Choose two from A-E in C(5,2)=10 ways, and from F-G in C(2,2) = 1 way.  That's
10 ways. 

4.	How many words can be made from the word “KEYBOARD” with any length?

1-letter words = P(8,1) = 8
2-letter words = P(8,2) = 8·7 = 56
3-letter words = P(8,3) = 8·7·6 = 336
4-letter words = P(8,4) = 8·7·6·5 = 1680
5-letter words = P(8,5) = 8·7·6·5·4 = 6720
6-letter words = P(8,6) = 8·7·6·5·4·3 = 20160
7-letter words = P(8,7) = 8·7·6·5·4·3·2 = 40320
8-letter words = P(8,8) = 8·7·6·5·4·3·2·1 = 40320

Add those up = 109600

Edwin